SHIFT CHECK LIST (you can find me in ∼cdaq/documents/FORMS/chklist e93026.ps)

DATE (month-day-year) Time (24 hour clock) Shift Checklist Recorder

Beam Energy (overview) (MeV) Beam Current (overview)(nA) Duty Factor (overview) Evidence of Scraping? (scope) HMS Trigger Rate (vis. scalers) Dead Time (vis. scalers) RunControl up? IPM3H00B nominal? (medm) Fast Raster Rmax,Ix,Iy (medm) Slow Raster OFF? (medm)

Run Number (recent one!) In /scalers/gen####.txt HMS WC fiducial effi. Computer Deadtime HMS Electronic Deadtime Charge scalers counting?

Electronics Room: WC Thresholds HMS/SOS BCM UP Temp. ok? HMS Trigger (eg., 3/4*anything) Electronics? (eg, unplugged, etc)

GAS SHED KEY is #29 Argon Bottle High/Low (PSI) Ethane Bottle High/Low (PSI) Helium Bottle High (PSI) Helium Flow (CFM) HMS Flow Ar CH3/Eth CH4 HMS Fridge (degC) Alcohol Bubbling? Recorded in Gas Logbook?

Column  1  2  3  4  5

Column  1  2  3  4  5

HMS/SOS Slits (LAT HCC18) Scattering Chamber Vacuum Target (material, length) HMS Shutter Open? (TV) HMS angle (TV/GUI) Hall Survey (TV) HMS Cer. Temp/Press (TV)

HMS Hodoscope HV HMS Drift Chamber HV HMS Gas Cerenkov HV HMS Pb-Glass HV

HMS Q1 Current polarity/setpt HMS Q1 Hall Field HMS Q2 Current polarity/setpt HMS Q2 Hall Field HMS Q3 Current polarity/setpt HMS Q3 Hall Field HMS D Current polar./readback HMS D NMR Field HMS Q1 N2/He Level HMS Q2 N2/He Level HMS Q3 N2/He Level HMS D N2/He Level

Neutron Counter HV(on/off) Threshold(mV) Neutron bars counter(10)(ok?) VPaddles count(6)(ok?)

hclog the following screens: BPM medm screen

scan into hclog when page is full Check List Complete!